The Bilt-Rite Electronics Company makes televisions, CD players, and DVD players. Each month, the company turns out 1750 units (TVs, CDs, and DVDs combined). They make twice the number of CDs as they make TVs. The number of DVDs they make is 100 more than the number of TVs and CDs combined. How many of each unit (TVs, CDs, and DVDs) does Bilt-Rite make each month?
I have set the unknowns as t, c, and d and I know that all together equals 1750, but after that I get very confused. Is it even possible to solve this?

Answer by ikleyn(52781)

.

From the condition, you have these equations  (I write them as I read)


    t + c + d = 1750     (1)

    c = 2t,              (2)

    d = t + c + 100.     (3)


Substitute expression (2) into equation (2). You get then


    d = t + 2t + 100 = 3t + 100    (4)


Now substitute (2) and (3) into (1) to get


    t + 2t + (3t+100) = 1750,

or

    6t                = 1750 - 100 = 1650.


Hence,  t = 1650/6 = 275.


From this point, complete the solution on your own.

Solved.